# StableSketcher: Enhancing Diffusion Model for Pixel-based Sketch Generation via Visual Question Answering Feedback

### v1

@@ Abstract
Although recent advancements in diffusion models have significantly enriched the quality of generated images, challenges remain in synthesizing pixel-based human-drawn sketches, a representative example of abstract expression. To combat these challenges, we propose StableSketcher, a novel framework that empowers diffusion models to generate hand-drawn sketches with high prompt fidelity. Within this framework, we fine-tune the variational autoencoder to optimize latent decoding, enabling it to better capture the characteristics of sketches. In parallel, we integrate a new reward function for reinforcement learning based on visual question answering, which improves text-image alignment and semantic consistency. Extensive experiments demonstrate that StableSketcher generates sketches with improved stylistic fidelity, achieving better alignment with prompts compared to the Stable Diffusion baseline. Additionally, we introduce SketchDUO, to the best of our knowledge, the first dataset comprising instance-level sketches paired with captions and question-answer pairs, thereby addressing the limitations of existing datasets that rely on image-label pairs. Our code and dataset will be made publicly available upon acceptance.
footnotetext: © 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This work has been submitted to IEEE Access for possible publication.
@@ 1 Introduction
The application of sketches spans diverse domains, including sketch-guided text-to-image generation [32, 39, 28], sketch-guided image editing [20, 35], and image retrieval [14, 25, 3], underscoring their significance in both creative and practical contexts. However, generative models often fail to capture the essence of human-drawn sketches, instead generating hyper-realistic renderings that deviate from the simplicity and abstraction inherent in sketches. Moreover, these models struggle with maintaining prompt fidelity, as illustrated in Figure 1.
To address these challenges, we propose StableSketcher, a framework that enhances the generative performance of Stable Diffusion [23] for abstract, human-drawn sketches. We fine-tune the variational autoencoder (VAE) of Stable Diffusion to optimize latent representations, ensuring stylistic coherence in generated outputs. Additionally, we define a novel reward function based on visual question-answering (VQA) feedback, integrating it into a reinforcement learning (RL) algorithm to improve the prompt fidelity of the generated sketches. Qualitative and quantitative evaluations, along with user studies, demonstrate that our framework outperforms the Stable Diffusion baseline in generating abstract sketches with improved prompt fidelity.
Along with the outlined issues, the development of robust sketch generation models has been hindered by the inherent limitations of existing sketch datasets [4, 5, 27, 21]. These datasets lack the semantic depth required for generative tasks, making them fit for sketch classification, but insufficient for text-to-image tasks. Furthermore, existing resources lack the fine-grained, instance-centric sketch–caption pairs required for sketch generation. Existing caption datasets [2, 22] describe relationships across multiple objects in a scene rather than the instance itself.
To combat these limitations, we propose SketchDUO, a comprehensive dataset containing 35.8K instance-level sketches paired with fine-grained textual captions and 54.3K question-answer (QA) pairs, offering rich semantic detail for modeling single-object sketches. SketchDUO includes both positive examples, reflecting the desired sketch style, and negative examples that capture common misrepresentations observed in Stable Diffusion outputs, such as sketches with excessive detail or shading. By incorporating contrastive examples, SketchDUO enhances the model’s understanding of desired and undesired styles, enabling it to generate sketches that better align with the intended style and fidelity.
@@ 2.1 Sketch Generation with Diffusion Models
Diffusion models are typically trained on large datasets of photorealistic images, resulting in a bias towards generating realistic, highly detailed outputs. This training bias limits their ability to generate abstract representations, such as sketches [26, 37, 15]. Furthermore, conventional diffusion models often lack fine-grained control over the structural and abstract elements in sketches, making it difficult to achieve the desired level of simplicity and abstraction [15, 33].
At the same time, the majority of sketch generation research has focused on vector-based and stroke-based approaches [5, 31], which capture sketches at the granularity of individual strokes. While such methods offer computational efficiency, they struggle to handle more complex and detailed sketches, often failing to capture the essence of human-drawn art. More recent work has pivoted to pixel-based generation using diffusion models. For instance,  [9] proposed a scale-adaptive diffusion model for sketch generation, employing a multi-step sampling technique to enhance the quality of the generated sketches. However, their method is constrained by the use of image-caption pairs, limiting its ability to effectively capture the desired style and characteristics of sketches.
Table 1: Comparison of sketch datasets. Prior sets target recognition; SketchDUO uniquely provides both captions and QA pairs.
Dataset # Classes # Sketches/ Class Total # Sketches Caption QA
TU-Berlin [4] 250 80 20K
Sketchy [27] 125 avg. 600 75K
QuickDraw [5] 345 avg. 144K $\sim$50M
SEVA [21] 128 avg. 703 90K
SketchDUO (ours) 30 avg. 1.2k 35.8K
@@ 2.2 Sketch Datasets
As research in generative models progresses, a growing variety of sketch datasets has emerged to support advancements in sketch-related studies. Table 1 compares several existing sketch datasets. QuickDraw [5] is one of the largest datasets for sketch classification, but the absence of annotations and low-quality sketches limit its applicability to generative tasks. TU-Berlin [4] provides more complex sketches, but it also lacks a description of instances. Sketchy [27] pairs sketches with images, but the complexity of the sketches and the reliance on image-label pairs limit its usability for sketch generation. More recent efforts, such as SEVA [21], use CLIPasso [31] to generate stroke-based sketches, yet these datasets still rely on photo-sketch pairs and lack the fine-grained, instance-level captions required for high-quality sketch generation.
@@ 2.3 Reinforcement Learning in Diffusion Models
The integration of reinforcement learning (RL) with generative models has garnered significant attention in recent years, particularly in the context of text-to-image generation. Reward-weighted maximum-likelihood estimation (RWR) has been widely used to align generated images with textual prompts [16]. However, methods that rely heavily on rewards can suffer from learning instability. To address this issue, denoising diffusion policy optimization (DDPO) [1] introduced a policy gradient-based reinforcement learning method that optimizes diffusion models directly, enabling users to define custom reward functions tailored to specific generative tasks. In DDPO, BERTScore [40] is used as a text-image alignment reward, but this approach is limited by the methodology that BERTScore relies on image captioning models, which hinders stability when representing abstract forms like sketches.
We leverage the TIFA score [10], which employs a question-answer set-based evaluation method to facilitate a more fine-grained assessment of text-image alignment. Expanding upon the TIFA score, we devise a novel VQA-based reward function, which is incorporated as feedback during training.
@@ 3 SketchDUO
SketchDUO contains 35,851 instance-level sketch images paired with textual captions and 54,370 question-answer (QA) sets. By offering both captions and QA pairs, SketchDUO provides rich descriptions for individual objects, effectively addressing the limitations of existing datasets. The dataset adopts a contrastive approach, featuring positive examples that capture the desired sketch style, and negative examples that highlight common misrepresentations in Stable Diffusion outputs, such as images with excessive detail, over-shading, or sketches that resemble photographs of pencil drawings rather than true hand-drawn representations.
@@ 3.1 Definition of a Sketch
Table 2: Examples from the SketchDUO dataset, showcasing both positive and negative sketches along with their corresponding captions and question–answer pairs.
Category Positive File Name fish_49.png
[Uncaptioned image] Caption A simple drawing of a fish with three curved lines on its body and a round eye on a white background.
Instance Q&A
$Q_{1}$: What animal is in the picture?
$A_{1}$: Fish
$Q_{2}$: How many lines are on the fish?
$A_{2}$: 3
Sketch Q&A
$Q_{1}$: Is the background white?
$A_{1}$: Yes
$Q_{2}$: Is this a simple or a complex drawing?
$A_{2}$: Simple
Category Negative File Name fish_2.png
[Uncaptioned image] Caption A detailed drawing of a blue and red fish with orange accents on a beige background featuring a lot of shading.
Instance Q&A
$Q_{1}$: What color is the fish?
$A_{1}$: Blue and red
$Q_{2}$: Are there orange accents on the fish?
$A_{2}$: Yes
Sketch Q&A
$Q_{1}$: Is there a detailed drawing?
$A_{1}$: Yes
$Q_{2}$: Is there a lot of or a little shading?
$A_{2}$: A lot of
Table 2 presents an example of images, captions, and QA sets from SketchDUO. To construct SketchDUO, we selected 30 common classes shared between the QuickDraw [5] and TU-Berlin [4] datasets. The selection of classes was designed to achieve a balanced representation of diverse objects, ensuring broad thematic coverage across the dataset.
SketchDUO comprises 30 classes, distributed across six broad categories to ensure balanced representation and thematic diversity. Below, we outline the six main categories, their corresponding classes, and their relative distributions:
@@ 3.2 Sketch Image Collection
Figure 3: Background-preserving sketch augmentations: original; rotation by $15^{\circ}$ with white padding; rotation by $15^{\circ}$ with white padding; Gaussian blur (weak, $k=3$, $\sigma=0.8$); Gaussian noise applied to strokes only (weak, $\sigma=8$); Gaussian blur (strong, $k=5$, $\sigma=1.6$); Gaussian noise applied to strokes only (strong, $\sigma=16$); and line thickening obtained by binarization followed by morphological dilation with a $3\times 3$ elliptical kernel for one iteration.
We curate a corpus of $24,000$ positive and $11,851$ negative sketch images. The positive portion is derived from $3,000$ human-drawn sketches spanning 30 classes, with one hundred sketches per class. The negative portion contains $1,693$ images generated with Stable Diffusion v2.1 and selected for off style traits such as intricate line work, the presence of color, nonwhite backgrounds, and heavy shading. To every image we apply the same seven background-preserving augmentations. For negative samples, which contain no strokes, we omit the line thickening augmentation. These include rotations of plus or minus fifteen degrees with white padding; Gaussian blur with a weak setting where $k$ equals three and sigma equals $0.8$, and a strong setting where $k$ equals five and $\sigma$ equals 1.6; Gaussian noise applied to strokes only with a weak setting where $\sigma$ equals eight and a strong setting where $\sigma$ equals sixteen; and line thickening achieved through morphological dilation, as detailed in Figure 3. The resulting collection contains $35,851$ images in total, comprising $24,000$ positive and $11,851$ negative samples. A comprehensive analysis of the distributions at the category and class levels for both subsets, including relative proportions and sample counts, is presented in Figure 2.
@@ 3.2.1 Positive Sketch Data Construction
To construct the positive samples, we collect 3,000 hand-drawn sketch images, with 100 instances per class. Each image was created by human participants following a standardized protocol to ensure stylistic and semantic consistency across the dataset. Participants were provided with reference examples that described the desired visual characteristics of positive sketches. All sketches were drawn using a tablet and Microsoft Paint (mspaint), the default drawing tool on Windows, to maintain uniformity in drawing tools and procedures. The images were saved in PNG format with a fixed resolution of $\times512$ pixels. The drawing style was carefully controlled to reflect the core properties of the dataset. Each sketch was required to depict a single object instance in a minimal style using thick black lines (#000000) on a pure white background (#FFFFFF). Drawings were intended to be completed within 40 to 60 seconds to encourage simplicity and consistency in visual abstraction. Captions for each image were initially generated using GPT-4o, and subsequently refined through human correction to improve alignment with the visual content. These finalized captions were then passed to a question generation module adapted from the TIFA framework [10], which leverages LLaMA 2 [30] for generating diverse questions and UnifiedQA [12] for answer validation. This process yielded structured QA sets for each image, ensuring that every positive sample in the dataset is paired with high-quality captions and reliable question–answer triplets.

### v2

@@ Abstract
Although recent advancements in diffusion models have significantly enriched the quality of generated images, challenges remain in synthesizing pixel-based human-drawn sketches, a representative example of abstract expression. To combat these challenges, we propose StableSketcher, a novel framework that empowers diffusion models to generate hand-drawn sketches with high prompt fidelity. Within this framework, we fine-tune the variational autoencoder to optimize latent decoding, enabling it to better capture the characteristics of sketches. In parallel, we integrate a new reward function for reinforcement learning based on visual question answering, which improves text-image alignment and semantic consistency. Extensive experiments demonstrate that StableSketcher generates sketches with improved stylistic fidelity, achieving better alignment with prompts compared to the Stable Diffusion baseline. Additionally, we introduce SketchDUO, to the best of our knowledge, the first dataset comprising instance-level sketches paired with captions and question-answer pairs, thereby addressing the limitations of existing datasets that rely on image-label pairs. Project page: https://zihos.github.io/StableSketcher
footnotetext: * Corresponding author.
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This work has been submitted to IEEE Access for possible publication.
@@ 1 Introduction
Sketches are particularly useful in scenarios that require rapid visual exploration and efficient communication of ideas. Such scenarios include early-stage concept ideation [33, 50, 28] and human–AI co-creative drawing [56, 23, 8], where simplified visual forms can facilitate iterative exploration and collaboration. The application of sketches spans diverse domains, including sketch-guided text-to-image generation [47, 59, 43, 5], sketch-guided image editing [32, 52, 30], and image retrieval [40, 7, 21, 41], underscoring their significance in both creative and practical contexts. However, generative models often fail to capture the essence of human-drawn sketches, instead generating hyper-realistic renderings that deviate from the simplicity and abstraction inherent in sketches. Moreover, these models struggle with maintaining prompt fidelity, as illustrated in Figure 1.
To address these challenges, we propose StableSketcher, a framework that enhances the generative performance of Stable Diffusion [38] for abstract, human-drawn sketches. We fine-tune the variational autoencoder (VAE) of Stable Diffusion to optimize latent representations, ensuring stylistic coherence in generated outputs. Additionally, we define a novel reward function based on visual question-answering (VQA) feedback, integrating it into a reinforcement learning (RL) algorithm to improve the prompt fidelity of the generated sketches. Qualitative and quantitative evaluations, along with user studies, demonstrate that our framework outperforms the Stable Diffusion baseline in generating abstract sketches with improved prompt fidelity. Along with the outlined issues, the development of robust sketch generation models has been hindered by the inherent limitations of existing sketch datasets [9, 11, 42, 34]. These datasets lack the semantic depth required for generative tasks, making them fit for sketch classification, but insufficient for text-to-image tasks. Furthermore, existing resources lack the fine-grained, instance-centric sketch–caption pairs. Existing caption datasets [6, 37] describe relationships across multiple objects in a scene rather than the instance itself.
To combat these limitations, we propose SketchDUO, a comprehensive dataset containing 35.8K instance-level sketches paired with fine-grained textual captions and 54.3K question-answer (QA) pairs, offering rich semantic detail for modeling single-object sketches. SketchDUO includes both positive examples, reflecting the desired sketch style, and negative examples that capture common misrepresentations observed in Stable Diffusion outputs, such as sketches with excessive detail or shading. The negative examples were introduced primarily as controlled counterexamples intended to expose traits the model should avoid, rather than to fully represent realistic human drawing errors. By incorporating contrastive examples, SketchDUO enhances the model’s understanding of desired and undesired styles, enabling it to generate sketches that better align with the intended style and fidelity.
@@ 2 Literature Review
Figure 2: SketchDUO construction process. The pipeline proceeds from left to right. It begins with class selection and sketch definition, followed by positive sketch collection and negative sketch construction. The two subsets are then augmented. In the caption generation stage, GPT-4o produces template-constrained instance-level captions for positive and negative sketches, and these captions are manually refined before TIFA-based QA pairs are constructed from each sketch–caption pair. The final SketchDUO dataset contains 35,851 sketch–caption pairs and 54,370 QA pairs.
Table 1: Representative examples from SketchDUO, showing positive–negative pairs for two object classes: fish (top two rows) and cake (bottom two rows).
[Uncaptioned image] Data Type positive File name fish_49.png
Caption A simple drawing of a fish with three curved lines on its body and a round eye on a white background.
Instance-related Question–Answer Q. What animal is in the picture? A. Fish Q. How many lines are on the fish? A. 3
Sketch-related Question–Answer Q. Is the background white? A. Yes Q. Is this a simple or a complex drawing? A. Simple
[Uncaptioned image] Data Type negative File name fish_2.png
Caption A detailed drawing of a blue and red fish with orange accents on a beige background featuring a lot of shading.
Instance-related Question–Answer Q. What color is the fish? A. Blue and red Q. How many lines are on the fish? A. 3
Sketch-related Question–Answer Q. Is there a detailed drawing? A. Yes Q. Is there a lot of or a little shading? A. A lot of
[Uncaptioned image] Data Type positive File name cake_26.png
Caption A simple drawing of a three-tier cake with three candles on a white background.
Instance-related Question–Answer Q. What is on the cake? A. Candles Q. Is the cake three-tier or two-tier? A. Three-tier
Sketch-related Question–Answer Q. What color is the background? A. White Q. Is this a simple or a complex drawing? A. Simple
[Uncaptioned image] Data Type negative File name cake_47.png
Caption A layered cake with alternating dark and light layers on a beige background featuring a lot of shading.
Instance-related Question–Answer Q. What type of food is this? A. Layered cake Q. What color are the layers? A. Dark and light
Sketch-related Question–Answer Q. What is the background color? A. Beige Q. Is there a lot of shading? A. Yes
Figure 4: Illustration of background-preserving sketch augmentations. The top row, from left to right, shows the original sketch, rotation by $15^{\circ}$ with white padding, rotation by $15^{\circ}$ with white padding, and weak Gaussian blur ($k=3$, $\sigma=0.8$). The bottom row presents weak Gaussian noise applied only to the stroke regions ($\sigma=8$), strong Gaussian blur ($k=5$, $\sigma=1.6$), strong Gaussian noise applied to the stroke regions ($\sigma=16$), and line thickening obtained through binarization followed by morphological dilation with a $3\times 3$ elliptical kernel for one iteration.
@@ 2.1 Sketch Generation with Diffusion Models
Diffusion models are typically trained on large datasets of photorealistic images, resulting in a bias towards generating realistic, highly detailed outputs. This training bias limits their ability to generate abstract representations, such as sketches [55, 27, 54, 22]. Furthermore, conventional diffusion models often lack fine-grained control over the structural and abstract elements in sketches, making it difficult to achieve the desired level of simplicity and abstraction [22, 49]. Although strategies to bridge the specific domain gap between sketches and photos have been extensively explored in heterogeneous face recognition through normalization and disentanglement frameworks [35, 36], effectively adapting these insights to the generative domain remains an open challenge.
At the same time, much of the sketch generation literature has focused on vector- and stroke-based representations, which model sketches at the granularity of individual strokes [11, 46]. While these approaches offer computational efficiency and editability, they often struggle to capture more complex and detailed sketches in a natural human-drawn style. Recent diffusion-based research has explored sketch synthesis in multiple representation spaces. For example, SwiftSketch [3] proposes an image-conditioned diffusion model for vector sketch generation, enabling efficient synthesis by progressively denoising stroke control points. In parallel, Hu et al. [15] extend diffusion models to pixel-level sketch generation through a scale-adaptive sampling strategy for complex sketch synthesis.
More recent work has also studied controllable diffusion-based sketch generation using continuous sketch representations, such as unsigned distance fields, to improve structural clarity and controllability [55]. However, these methods either focus on image-conditioned vector synthesis or depend on representations and training settings that do not directly address pixel-based generation of human-drawn sketches with fine-grained textual fidelity.
@@ 2.2 Sketch Datasets
As research in generative models progresses, a growing variety of sketch datasets has emerged to support advancements in sketch-related studies. QuickDraw [11] is one of the largest datasets for sketch classification, but the absence of annotations and low-quality sketches limit its applicability to generative tasks. TU-Berlin [9] provides more complex sketches, but it also lacks descriptions of individual instances. Sketchy [42] pairs sketches with images, but the complexity of the sketches and the reliance on image-label pairs limit its usability for sketch generation. More recent efforts, such as SEVA [34], leverage CLIPasso [46] to construct stroke-based sketch data for evaluating abstraction, yet they still rely on photo-sketch pairs and do not provide the fine-grained, instance-level captions needed for high-quality text-to-sketch generation. Likewise, SketchRef [29] emphasizes benchmark construction and evaluation metrics for automated sketch synthesis, focusing on recognizability, structural consistency, and fairness across different simplification levels rather than providing richly annotated instance-level sketch–caption pairs for generative training. Therefore, existing datasets remain limited for learning pixel-based sketch generation with fine-grained semantic alignment.
@@ 2.3 Reinforcement Learning in Diffusion Models
The integration of reinforcement learning (RL) with diffusion models has recently attracted increasing attention, particularly for text-to-image alignment. Early work explored reward-weighted regression (RWR) to align generated images with textual prompts [24], but such reward-weighted objectives can suffer from instability and limited credit assignment. More direct RL formulations have therefore been proposed for diffusion fine-tuning. For example, diffusion policy optimization with KL regularization (DPOK) [10] formulates text-to-image diffusion fine-tuning as an online RL problem with policy gradient optimization and KL regularization. Building on this perspective, denoising diffusion policy optimization (DDPO) [4] models the denoising process itself as a sequential decision-making problem, allowing diffusion models to be optimized directly with user-defined reward functions. More recent work has further extended RL-based diffusion training to address sparse-reward settings during denoising [17].
DDPO is particularly relevant to our setting because it enables direct optimization with task-specific black-box rewards defined on generated images. In the original DDPO framework, BERTScore [58] is used as a text-image alignment reward, but such caption-based similarity is limited for abstract forms like sketches. We instead leverage the TIFA score [16], which uses question-answer based evaluation for a more fine-grained assessment of text-image alignment. Building on TIFA, we devise a novel VQA-based reward function and incorporate it into DDPO training to improve the semantic fidelity of generated sketches.
@@ 3 SketchDUO
SketchDUO contains 35,851 instance-level sketch images paired with textual captions and 54,370 QA sets. By offering both captions and QA pairs, SketchDUO provides rich descriptions for individual objects, effectively addressing the limitations of existing datasets. The dataset adopts a contrastive approach, featuring positive examples that capture the desired sketch style, and negative examples that highlight common misrepresentations in Stable Diffusion outputs, such as images with excessive detail, over-shading, or sketches that resemble photographs of pencil drawings rather than true hand-drawn representations. SketchDUO was constructed through a human-in-the-loop pipeline, in which model-generated captions and QA candidates were subsequently verified and revised by human annotators to ensure semantic accuracy and subset-specific stylistic consistency.
Figure 2 provides an overview of the SketchDUO construction pipeline, and Table 1 presents example images, captions, and QA sets from SketchDUO. To construct SketchDUO, we selected 30 common classes shared between the QuickDraw [11] and TU-Berlin [9] datasets. The selection of classes was designed to achieve a balanced representation of diverse objects, ensuring broad thematic coverage across the dataset. SketchDUO comprises 30 classes, distributed across six broad categories to ensure balanced representation and thematic diversity. Below, we outline the six main categories and their corresponding classes:
@@ 3.2 Sketch Image Collection
Table 2: GPT-4o captioning prompts and human revision criteria used in SketchDUO construction. The positive and negative datasets use different prompt templates to reflect the target sketch style and controlled off-style contrastive examples, respectively.
Positive dataset  Model: GPT-4o System message: You are a helpful assistant for generating image captions. Input: An image provided as a base64-encoded JPEG string. Output template “A black line drawing of {{text1}} on a white background.” or “A simple drawing of {{text1}} on a white background.” Key prompt constraints – Replace {{text1}} with a detailed instance-level description. – Focus on clear objects, shapes, and actions rather than vague category-only descriptions. – Choose “black line drawing” for relatively detailed sketches and “simple drawing” for more minimal sketches. – Do not output the placeholder brackets. Human revision criteria – Correct inaccurate object attributes, counts, and local structural details. – Remove vague wording. – Ensure consistency with the positive sketch definition: a single-object black-line drawing on a white background with minimal detail. Negative dataset  Model: GPT-4o System message: You are a helpful assistant for generating image captions. Input: An image provided as a base64-encoded JPEG string. Output template “A {{text1}} drawing of {{text2}} on a {{text3}} background {{text4}}.” Key prompt constraints – Replace {{text1}} with “detailed” or “simple” according to sketch complexity. – Replace {{text2}} with a detailed object description. – Replace {{text3}} with the background color. – Replace {{text4}} with “featuring a lot of shading.” when shading is present; otherwise leave it blank. – This template explicitly allows off-style traits such as color, non-white background, and shading. Human revision criteria – Correct semantic errors while preserving contrastive off-style properties. – Retain color, excessive detail, non-white background, and shading when visually present. – Preserve the negative subset’s role as controlled off-style supervision.
We curate a corpus of $24,000$ positive and $11,851$ negative sketch images. The positive portion is derived from $3,000$ human-drawn sketches spanning 30 classes, with one hundred sketches per class. The negative portion contains $1,693$ images generated with Stable Diffusion v2.1 and selected for off style traits such as intricate line work, the presence of color, nonwhite backgrounds, and heavy shading.
To every image we apply the same seven background-preserving augmentations. For negative samples, which contain no strokes, we omit the line thickening augmentation. These include rotations of plus or minus fifteen degrees with white padding; Gaussian blur with a weak setting where $k$ equals three and sigma equals $0.8$, and a strong setting where $k$ equals five and $\sigma$ equals 1.6; Gaussian noise applied to strokes only with a weak setting where $\sigma$ equals eight and a strong setting where $\sigma$ equals sixteen; and line thickening achieved through morphological dilation, as detailed in Figure 4. The resulting collection contains $35,851$ images in total, comprising $24,000$ positive and $11,851$ negative samples. A comprehensive analysis of the distributions at the category and class levels for both subsets, including relative proportions and sample counts, is presented in Figure 3.
@@ 3.2.1 Positive Sketch Data Construction
To construct the positive samples, we collected 3,000 hand-drawn sketch images, with 100 instances per class. These sketches were created by 12 volunteer participants, none of whom were art majors. All participants followed a standardized drawing protocol to ensure stylistic and semantic consistency across the dataset. They were provided with reference examples illustrating the desired visual characteristics of positive sketches. All sketches were drawn using a tablet and pen with the default drawing tool to maintain consistency in the drawing tools and procedures. The images were saved in PNG format at a fixed resolution of $\times512$ pixels. The drawing style was carefully controlled to reflect the core properties of the dataset. Each sketch was required to depict a single object instance in a minimal style using thick black lines (#000000) on a pure white background (#FFFFFF). Participants were instructed to complete each drawing within 40 to 60 seconds to encourage simplicity and consistency in visual abstraction.
Captions for each image were initially generated with GPT-4o using a template-constrained prompt tailored to the positive sketches. The prompt restricted the output to a single-object black-on-white sketch description, consistent with our sketch definition. Each generated caption was then manually revised to correct object attributes, counts, and local structural details before QA construction. The exact prompt is provided in Table 2.
These finalized captions were then passed to a question generation module adapted from the TIFA framework [16], which leverages LLaMA 2 [45] for generating diverse questions and UnifiedQA [19] for answer validation. This process yielded structured QA sets for each image, ensuring that every positive sample in the dataset is paired with high-quality captions and reliable question–answer triplets.